In the process of rubbing the lenses of the eyeglasses, 6.28x1010 electrons were transferred. 

a. What is the charge of the lenses and the nylon cloth?

b. What is the change in their masses?



1
Expert's answer
2021-02-28T07:37:50-0500

(a) The charge of the lenses

"q_1=eN\\
=1.6\times 10^{-19}\:\rm C\times 6.28\times 10^{10}=1.00\times 10^{-8}\:\rm C"

(b) The charge of the nylon cloth

"q_2=-eN\\
=-1.6\times 10^{-19}\:\rm C\times 6.28\times 10^{10}=-1.00\times 10^{-8}\:\rm C"